• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen/xiaohui

    Log entry time 13:50:22 on July23,2006

    Entry number 175007

    keyword=LH2/LD2 operation. LH2 Temp raised to 19.5K

    Loop 3 (LD2) JT was not responding. After JP came in, it worked.
    We adjusted LD2 JT to have HP of 160-170 W. With 20 uA on, it still have about 20-30 W. Also raised LH2 operating temperature from 19 K
    to 19.5 K (from run 1585), which gave us about 70 W extra HP power.
    We lowered loop 1 JT to keep LH2 (loop 1) to have about 140 W. When
    20 uA on, it has 20-30 W. The total coolant flow is about 15.7 g/s.
    So for the next a few hours of running, we will not need to adjust
    JT or fan speed, until Hall C needs more power (Hall C is using 6 g/s now).